OpenStudy (anonymous):

How do i find the smaller root to 6x^2=6 ?

OpenStudy (anonymous):

First move everything to one side: 6x^2 - 6 = 0 now factor out a 6: 6(x^2 - 1) = 0 6(x - 1) (x + 1) = 0 By definition a root is any x-value that makes the equation 0, in this case the roots would be: x = 1,-1 the smallest root therefore is -1

OpenStudy (anonymous):

hi julie the smalles root is -1